Please help.
the amount of light (measured in foot-candles) produced by a light source varies inverse;y as the square of the distance from the source. If the illumination produced 1 m from a light source is 768 foot-candles, find the illumination produced 6 m from the same source.

Answer by stanbon(75887)

the amount of light (measured in foot-candles) produced by a light source varies inverse;y as the square of the distance from the source. If the illumination produced 1 m from a light source is 768 foot-candles, find the illumination produced 6 m from the same source.
---------------------
L = k/d^2
Solve for "k"::
768 = k/1^2
k = 768
---------------
Equation:
L = 768/d^2
-----
Ans: L = 768/6^2 = 21 1/3 ft-candles
-------------------
Cheers,
Stan H.
-------------------